OpenStudy (anonymous):

what is 2/5 divided by 3 1/2

OpenStudy (anonymous):

\[\frac{\frac{2}{5}}{3.5}=\frac{2}{5}*\frac{1}{3.5}=\frac{2}{17.5}\]

OpenStudy (anonymous):

Or since \[3.5=\frac{7}{2}\]Then\[\frac{\frac{2}{5}}{\frac{7}{2}}=\frac{2}{5}*\frac{2}{7}=\frac{4}{35}\]